OpenStudy (anonymous):

Use a graph and a table to solve the system. Check your answer. Enter your answer as the following example: (5, -6) x-y=-1 4x-2y=2

OpenStudy (anonymous):

plot these two lines on a graph...

OpenStudy (anonymous):

1st make tables for each line and then plot each of them on d graph

OpenStudy (anonymous):

im guessing it's were they meet

OpenStudy (anonymous):

it'll meet at (2,3)

OpenStudy (anonymous):

x-y=-1 2x-y=1...now u can can tell..whether they'll meet or not?

OpenStudy (anonymous):

they meet at (0,1)

OpenStudy (anonymous):

Basically, you use the substitution method to find the values of both the y and x variable; the tables just allow you to keep organized. Graphing helps you visualize as well. So, start out by pushing numbers. Given: x-y=-1 4x-2y=2 substitute "0" for x: 0 - y = -1 x - (1) = -1 0 = -1 + y x - 1 = -1 1 = y x = 0 meeting point is thus : (0,1) To find yet another point, just use the point slope equation - after all, you can make a line with only two points